Question 2

Using the financial information below, calculate company XYZ's 2017 Liabilities to Shareholder Equity ratio. Enter your answer as a decimal with four places of precision (i.e. 0.1234).

xid-100948877_1
Answer #1

Liabilities to shareholders equity ratio = Total liabilities / Total Equity

From the given data Total liabilities = Account Payable+ Deferred Revenue+Net note payable

Total Equity = Common Stock + Retained Earnings

Total Liabilities = 5000+1000+ 10000-800 = 15200

Total Equity = 12,000+4,800 = 16,800

Ratio = 15,200 / 16,800 = .9047